OpenStudy (anonymous):

which equation represents the line parallel to the line y=2x-4

jimthompson5910 (jim_thompson5910):

Anything of the form y = 2x + b where b is any number will be parallel to y = 2x-4 the two lines, y = 2x+b and y = 2x-4, both have the same slope of m = 2

jimthompson5910 (jim_thompson5910):

for example, if b = 7, then y = 2x+b turns into y = 2x+7